# Reliability and Interpretability in Science and Deep Learning

### 14 Jan 2024 v1

@@ 4 Assumptions, simplicity and interpretability
Any model, whether TS or ML is essentially characterised by all its assumptions and its accuracy with respect to the existing empirical data [^10] Defining a scientific or ML model by its assumptions might suggest that our discussion is restricted to a syntactic approach to scientific models. However, following (Lutz, 2017, 2014), it also applies to a semantic approach. Other views exist in the philosophical literature (Winther, 2021). But, there is currently no compelling evidence of real models that do not admit a formal description.. In the following, the expression “all assumptions” refers to all those that are necessary to reproduce the outcome of the model, including the comparison with the empirical evidence and the expression of the input data [^11] If the model predicts an outcome with some statistical distribution, as it is mostly the case, the requirement is to reproduce the statistical distribution and not the individual outcomes.. The training data themselves are not part of the model assumptions [^12] More precisely, they do not matter in any comparison. See footnote 31.: they contribute to the empirical accuracy of the model just like the experiments that contribute to the development of TS models.
@@ 4.1 Assumptions
TS models rely on multiple categories of assumptions: besides the specific assumptions of the model itself, they rely also on multiple scientific disciplines that are not the main focus of the specific TS model but are essential to constrain the TS model. These disciplines are referred to as background science and they include a variety of domains from logic, mathematics, basic physics, chemistry and the modelling of any relevant experimental device. TS models typically include also assumptions to restrict their own domain of applicability [^13] Note that scope restrictions are logically equivalent to additional assumptions and are treated here as such.. Finally, TS models assume that their free parameters should be selected based on either a Maximum Likelihood principle (in a frequentist approach) or via a Bayesian argument that includes the assumption of Bayesian priors. Notably, TS models do not normally question the parameters inherited from background science, unless it becomes really compelling. In fact, those parameters must be consistent with a much broader spectrum of empirical evidence from very different domains. In this sense, traditional science adopts a kind of divide et impera strategy to parameter fitting.
On the other hand, DNN models require very little assumptions from background science (besides modelling the data collection). They are very generic and they are essentially defined only by the architecture parameters and hyper-parameters. However, their predictions are not determined by the training data, the DNN architecture and the Maximum Likelihood principle alone, not even approximately. They may depend also on subtle details of the training process. In fact, weight initialisation and pre-training techniques are key design choices when training and deploying a DNN model (Narkhede et al., 2022; Glorot and Bengio, 2010). Adversarials (Szegedy et al., 2014) are also evidence of high sensitivity to details. If we had evidence that the training of a DNN model depends only on some simple rules and is independent of the specific initial values of the weights (within some clearly defined bound and approximations), we would need to include only these rules as part of the assumptions of the DNN model. Instead, the lack of understanding of the training dynamics enforces the inclusion of the full specification of the DNN (initial) parameters as part of the assumptions.
@@
Another interesting comparison is with RF models: they also display significant sensitivity to some details (initial conditions and algorithmic hyperparameters), but they are based on limited and well-defined features. This enables, in the first place, a clear definition of the domain of applicability. It also makes it easier to identify conditions of robustness, for some applications and contexts. A deep analysis of the stability of RF models is beyond the scope of this article. However, it should be noted that their sensitivity to details seems closer to traditional applications of MCMC rather than DNN. Tab. 1 briefly summarises the content of this section.
Assumptions DNN RF LR TS
(a) background Little/None Features Features Many domains
(b) domain Data collection set-up Features Features Features
(c) model Network Architecture Features, hyper-param Features Many
(d) parameters Init. weights & biases Init. params MaxLH MaxLH / Priors
Table 1: A brief overview of the main categories of assumptions needed for DNN, RF, LR and TS models.

### 31 Jan 2024 v2

@@ 4 Assumptions, simplicity and interpretability
Any model, whether TS or ML is essentially characterised by all its assumptions and its accuracy with respect to the existing empirical data [^10] Defining a scientific or ML model by its assumptions might suggest that our discussion is restricted to a syntactic approach to scientific models. However, following (Lutz, 2017, 2014), it also applies to a semantic approach. Other views exist in the philosophical literature (Winther, 2021). But, there is currently no compelling evidence of real models that do not admit a formal description.. In the following, the expression “all assumptions” refers to all those that are necessary to reproduce the outcome of the model, including the comparison with the empirical evidence and the expression of the input data [^11] If the model predicts an outcome with some statistical distribution, as it is mostly the case, the requirement is to reproduce the statistical distribution and not the individual outcomes.. The training data themselves are not part of the model assumptions [^12] More precisely, input data are included, but they do not matter in any comparison. See 32.: they contribute to the empirical accuracy of the model just like the experiments that contribute to the development of TS models.
@@ 4.1 Assumptions
TS models rely on multiple categories of assumptions: besides the specific assumptions of the model itself, they rely also on multiple scientific disciplines that are not the main focus of the specific TS model but are essential to constrain the TS model. These disciplines are referred to as background science and they include a variety of domains from logic, mathematics, basic physics, chemistry and the modelling of any relevant experimental device. TS models typically include also assumptions to restrict their own domain of applicability [^13] Note that scope restrictions are logically equivalent to additional assumptions and are treated here as such.. In particular, TS models assume suitable values of the model parameters [^14] Alternatively, one might assume that the free parameters are determined by Maximum Likelihood principle (in a frequentist approach) or via a Bayesian argument that includes the assumption of Bayesian priors. However, this requires also the assumption of the all the empirical evidence necessary to compute the likelihood function, which is typically less concise than assuming the values of the parameters (see Sec. 4.2 and Footnote 32).. The parameters inherited from background science must be consistent with a much broader spectrum of empirical evidence from very different domains. In this sense, traditional science adopts a kind of divide et impera strategy to parameter fitting.
On the other hand, DNN models require very little assumptions from background science (besides modelling the data collection). They are very generic and they are defined, ideally, only by the architecture (hyper-)parameters and the domain in which they are trained. However, this is not precise enough to determine their predictions, not even statistically. Predictions are certainly determined by the entire set of DNN parameters, which are typically in the order of billions in modern DNNs. Ideally, the specific values of all those parameters are not essential to determine a prediction, which should depend only on the training domain. However, the training domain is difficult to define and the outcome may depend also on subtle details of the training process. In fact, weight initialisation and pre-training techniques are key design choices when training and deploying a DNN model (Narkhede et al., 2022; Glorot and Bengio, 2010). Adversarials (Szegedy et al., 2014) are also evidence of high sensitivity to details. If we had evidence that the training of a DNN model depends only on some simple rules and is independent of the specific initial values of the weights (within some clearly defined bound and approximations), we would need to include only these rules as part of the assumptions of the DNN model. Instead, the lack of understanding of the training dynamics enforces the inclusion of the full specification of the DNN (initial) parameters and training data as part of the assumptions.
@@
Another interesting comparison is with RF models: they also display significant sensitivity to some details (initial conditions and algorithmic hyperparameters), but they are based on limited and well-defined features. This enables, in the first place, a clear definition of the domain of applicability. It also makes it easier to identify conditions of robustness, for some applications and contexts. A deep analysis of the stability of RF models is beyond the scope of this article. However, it should be noted that their sensitivity to details seems closer to traditional applications of MCMC rather than DNN.